OpenStudy (anonymous):

company a charges $25 per month plus $0.22 per minute for cell phone service. company b charges $35 per month plus $0.12 per minute. how many mintues of usage is needed in one month to make company b a better deal ?

OpenStudy (anonymous):

25 + 0.22x = 35 + 0.12x 0.22x - 0.12x = 35-25 0.10x = 10 x = 10 / 0.10 x = 100 The answer is more than any more than 100 minutes makes company b a better deal than company a.